# Semi-Supervised Risk Control via Prediction-Powered Inference

### v1

@@ VII-B Early Time Series Classification
Figure 3 (a) illustrates the accumulated halt times (top) and empirical conditional accuracy gap (bottom), when using 300 labeled calibration samples. Following the top panel of that figure, the SS-RCPS CLT achieves the earliest halt times, but this method is not guaranteed to be valid in finite samples. The SS-RCPS WSR does not stop early at all when the sample size is small, which goes in hand with the ImageNet experiment. The SS-RCPS BIN method performs better than RCPS lab for most time steps, both in terms of earlier halt times and variance reduction. Figure 3 (b) repeats the same experiment except that we now use a larger labeled calibration set of size 4000 samples. Here, the SS-RCPS BIN method loses its advantage over the other methods, while SS-RCPS WSR outperforms RCPS lab and performs comparably to SS-RCPS CLT. This conclusion is consistent with the ImageNet experiment, where SS-RCPS WSR demonstrated better performance with larger sample sizes of the labeled calibration data. Notably, all the methods control the conditional accuracy gap, indicated by the bottom panels of Figures 3.
@@ VIII Conclusions
We presented a novel approach for effectively leveraging unlabeled data to improve the statistical efficiency of RCPS. The key properties of our methods are as follows: (1) when the imputation model is relatively accurate, our semi-supervised calibration procedures successfully reduce the conservatism and variance of the prediction rule’s error rate; and (2) all the presented methods provably remain valid even when handed inaccurate imputations. With that said, when the imputations are inaccurate our approach may not improve the statistical efficiency of the vanilla RCPS which relies only on labeled samples. This calls for the design of imputation strategies and score functions to mitigate this issue, as an alternative to using the most likely predicted label. Another limitation is the i.i.d. assumption required to ensure the validity of our methods. A promising future direction would be to extend our approach to handle covariate (or label) shift, possibly by a reweighting mechanism of the loss functions [34, 10, 24].
@@ Acknowledgments
This research was supported by the Israel Science Foundation (grant No. 729/21). Y. R. thanks the Career Advancement Fellowship, Technion, for providing research support. We would like to thank Anastasios N. Angelopoulos for insightful comments regarding the semi-supervised calibration procedure for a general loss.
@@ A-A Fixed Sequence Testing
We begin by presenting more details regarding RCPS when the risk is not monotonous with respect to the hyper-parameter $q$. In this case, the risk-control task is reformulated as a multiple-hypothesis testing task. We pre-specify a grid search for the parameter $\hat{q}$: $\mathcal{Q}_{\text{grid}}=\{q_{1},q_{2},\ldots,q_{N}\}$ and compute a UCB with error level $\delta$ at these specifies points $(\hat{R}^{+}(q_{1}),\hat{R}^{+}(q_{1}),\ldots,\hat{R}^{+}(q_{N}))$. For each value, we define the null hypothesis: $H_{0,i}:R(q_{i})>\alpha,$ indicating that the risk is not controlled. We iterate over the grid and set $\hat{q}$ to be $q_{i}$ if $\hat{R}(q_{i})\leq\alpha$. We stop the iteration at the first instance where we cannot reject the null hypothesis. This procedure is called fixed-sequence-testing and is depicted in the following algorithm.
@@ B-B Variance Reduction
Denote by $L,\tilde{L}$ two random variables that take values in $\{0,1\}$. Further denote by $p_{1},p_{2}$ the probabilities: $\mathbb{P}(L=1,\tilde{L}=0)$ and $\mathbb{P}(L=0,\tilde{L}=1)$, respectively. In what follows, we show that
@@ Appendix C Additional Experimental Details
C-A Score Definition: Classification Experiment

### v2

@@ VII-B Early Time Series Classification
Figure 3 (a) illustrates the accumulated halt times (top) and empirical conditional accuracy gap (bottom), when using 300 labeled calibration samples. Following the top panel of that figure, the SS-RCPS CLT achieves the earliest halt times, but this method is not guaranteed to be valid in finite samples. The SS-RCPS WSR does not stop early at all when the sample size is small, which goes in hand with the ImageNet experiment. The SS-RCPS BIN method outperforms RCPS lab for most time steps, both in terms of earlier halt times and variance reduction. Figure 3 (b) repeats the same experiment except that we use a larger labeled calibration set of size 4000 samples. Here, the SS-RCPS BIN method loses its advantage over the other methods, while SS-RCPS WSR outperforms RCPS lab and performs comparably to SS-RCPS CLT. This conclusion is consistent with the ImageNet experiment, where SS-RCPS WSR demonstrated better performance with larger labeled calibration data. Notably, all methods control the conditional accuracy gap, indicated by the bottom panels of Figures 3.
@@ VIII Conclusions
We presented a novel approach for effectively leveraging unlabeled data to improve the statistical efficiency of RCPS. The key properties of our methods are as follows: (1) when the imputation model is relatively accurate, our semi-supervised calibration procedures successfully reduce the conservatism and variance of the prediction rule’s error rate; and (2) all the presented methods provably remain valid even when handed inaccurate imputations. With that said, when the imputations are inaccurate our approach may not improve the statistical efficiency of the vanilla RCPS which relies only on labeled samples. This calls for the design of imputation strategies and score functions to mitigate this issue, as an alternative to using the most likely predicted label. Another limitation is the i.i.d. assumption required to ensure the validity of our methods. A promising future direction would be to extend our approach to handle covariate (or label) shift, possibly by a reweighting mechanism of the loss functions [34, 10, 24]. Another promising direction is to derive a p-value or UCB for a trinomial variable, which is the actual loss we aim to control, thereby eliminating the need for the conservative clipping operator.
@@ Acknowledgments
This research was supported by the European Union (ERC, SafetyBounds, 101163414). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. This research was also partially supported by the Israel Science Foundation (ISF grant 729/21). Y. R. acknowledges additional support from the Career Advancement Fellowship at the Technion. The authors are deeply grateful to Anastasios N. Angelopoulos for his insightful comments and valuable feedback.
@@ A-A Fixed Sequence Testing
We begin by presenting more details regarding RCPS when the risk is not monotonous with respect to the hyper-parameter $q$. In this case, the risk-control task is reformulated as a multiple-hypothesis testing task. We pre-specify a grid search for the parameter $\hat{q}$: $\mathcal{Q}_{\text{grid}}=\{q_{1},q_{2},\ldots,q_{N}\}$ and compute a UCB with error level $\delta$ at these specifies points $(\hat{R}^{+}(q_{1}),\hat{R}^{+}(q_{2}),\ldots,\hat{R}^{+}(q_{N}))$. For each value, we define the null hypothesis: $H_{0,i}:R(q_{i})>\alpha,$ indicating that the risk is not controlled. We iterate over the grid and set $\hat{q}$ to be $q_{i}$ if $\hat{R}(q_{i})\leq\alpha$. We stop the iteration at the first instance where we cannot reject the null hypothesis. This procedure is called fixed-sequence-testing and is depicted in the following algorithm.
@@ A-B Waudby-Smith-Ramdas UCB
[…]
Figure 4: ImageNet classification experiment: the coverage rate of the constructed prediction sets as a function of the number of labeled calibration samples, evaluated over 100 different random splits of the data. Comparison between the modified WSR algorithm for bounded random variables and the original WSR using linear scaling. Other details are as in Figure 1.
@@ B-B Variance Reduction
[…]
@@ Appendix C Additional Experimental Details
C-A Classification Experiment
C-A1 Score Definition
@@
C-A2 Power Tuning
In this section, we provide further details on the power parameter $\lambda$ from [43] (introduced in Section V) and two variants of our method that incorporate it to attain more controlled variance reduction.
In more detail, the authors of [43] suggest a variant of the prediction-powered risk from (6) by introducing a hyperparameter $\lambda$ that controls the influence of the unlabeled data:
$\hat{R}_{\text{PP}}^{++}(q)=$ (37)
$\lambda\frac{1}{N}\sum_{j=n+1}^{n+N}\tilde{L}_{j}(q)+\frac{1}{n}\sum_{i=1}^{n}L_{i}(q)-\lambda\frac{1}{n}\sum_{i=1}^{n}\tilde{L}_{i}(q).$ (38)
Notice that the above risk remains unbiased for a fixed $\lambda$. The optimal parameter $\lambda^{*}$ is then obtained by minimizing the variance of this risk, yielding:
$\lambda^{*}=\frac{\text{Cov}(L_{i}(q),\tilde{L}_{i}(q))}{(1+\frac{n}{N})\text{Var}(\tilde{L}_{i}(q))}.$
Since $\text{Cov}(L_{i}(q),\tilde{L}_{i}(q))$ and $\text{Var}(\tilde{L}_{i}(q))$ are unknown, these must be estimated from the data. As a result, the parameter $\lambda^{*}$ aims to tighten the semi-supervised UCBs, ensuring they outperform or match those computed with labeled data alone. Notably, since $\text{Cov}(L_{i}(q),\tilde{L}_{i}(q))$ and $\text{Var}(\tilde{L}_{i}(q))$ are estimated on the same holdout data used to construct the UCB, this approach is valid only asymptotically [43]. Conversely, the construction of a finite-sample valid UCB requires additional hold-out data—or splitting the limited available calibration data—to estimate $\lambda^{*}$.
We now turn to introduce two additional variants of our method that utilize $\lambda$ for power tuning. The first computes the UCB using the CLT, as in Algorithm 1, but for the modified risk $\hat{R}_{\text{PP}}^{++}(q)$ from (37). In this case, the guarantee is only asymptotic, and thus we can estimate $\lambda^{*}$ with the same hold-out data used for computing the UCB. Figure 5 compares the performance of RCPS with labeled data, the original CLT-based UCB with $\lambda=1$ (as in Figure 1), and the CLT-based UCB with the estimated $\lambda^{*}$. As can be seen, the latter indeed yields the least conservative prediction sets, as expected.
Figure 5: ImageNet classification experiment: the coverage rate of the constructed prediction sets, as a function of the number of labeled calibration samples evaluated over 100 different random splits of the data. Comparison between RCPS using only labeled data, the CLT UCB with $\lambda=1$, and the CLT UCB with the tuned $\lambda$. Other details are as in Figure 1.
The second variant is the WSR-based UCB from Algorithm 1, which is supported by finite sample guarantees. Since the focus here is on non-asymptotic validity, estimating $\lambda^{*}$ requires additional data splitting. With limited labeled data, this is not desirable, as it makes the UCB more conservative due to the smaller sample size. However, with a relatively large amount of data, this method can be beneficial. Figure 6 compares RCPS using labeled data only, the WSR UCB with $\lambda=1$, and the WSR UCB with data splitting used to estimate $\lambda^{*}$. The setup uses 2000 labeled holdout points (200 for estimating $\lambda^{*}$ and the rest for UCB construction), and 23,000 unlabeled points (2300 for $\lambda^{*}$ estimation, and the remainder for UCB construction). Following that figure, we can see that for lower accuracies, the use of the estimated $\lambda^{*}$ is indeed beneficial: the estimated $\lambda^{*}$ effectively balances the reliance on unlabeled data. However, for higher accuracies, splitting the data slightly reduces statistical efficiency. This is because the optimal $\lambda^{*}$ is approximately 111 as in the WSR UCB method from the main manuscript that uses a fixed $\lambda=1$, but the UCB of the tuned method is now constructed with fewer samples.
Figure 6: ImageNet classification experiment: the coverage rate of the constructed prediction sets as a function of the imputation accuracy, evaluated over 100 different random splits of the data. Comparison between RCPS using only labeled data, the original WSR UCB that uses $\lambda=1$, and the WSR UCB with the tuned $\lambda^{*}$. Other details are as in Figure 2.